# IndiSeek learns information-guided disentangled representations

### v1

@@ 5 Experiments with MultiBench datasets
In addition to the application in single-cell multi-omics, we also evaluate the performance of IndiSeek on MultiBench datasets (Liang et al.,, 2021), which include 4 video datasets (MOSI (Zadeh et al.,, 2016), MOSEI (Zadeh et al.,, 2018), UR-FUNNY (Hasan et al.,, 2019), MUStARD (Castro et al.,, 2019)) and Medical Information Mart for Intensive Care III (MIMIC) dataset (Johnson et al.,, 2016). We follow the processing pipeline in Liang et al., (2023) and adopt exactly the same data splitting and feature pre-extraction, in which we focus on the task of binary classification on whether the patient fits any ICD-9 code in group 7 for the MIMIC dataset. We use a smaller Transformer architecture with $2$ heads, $2$ layers, and the intermediate layers with a width $128$. For all methods in comparison, we train the architecture for $2000$ epochs with the same training parameters. We note that our focus here is to benchmark the qualities of representations learned with task-agnostic objective functions. Therefore, for fair comparison along this direction, we do not involve task-related within- or cross-modality data augmentations and hence do not involve self-supervised learning terms in objective functions of all methods in comparison. In contrast, to achieve the optimal performance in a specific task on a particular benchmark dataset, appropriate data augmentation informed by the task and domain knowledge is often helpful.
Following Liang et al., (2023) and Wang et al., 2024a , we focus on the linear probing accuracy for each method with learned representations $(C_{1},Z_{1},C_{2},Z_{2})$ as features, i.e., we concatenate shared and modality-specific features for classification. Here we also compare with the CLIP baseline, where only the shared representations $(C_{1},C_{2})$ are used for linear probing accuracy evaluation. Following the same training procedure as for the CITE-seq dataset, we report the results with three methods in Table 1. For each method, we vary $\lambda$ in $\{10^{j}:j=-3,-2,-1,0,1,2,3\}$ and for each $\lambda$, we run each method with $10$ random seeds. Table 1 presents the averaged accuracy for the best choice of $\lambda$. More details on the performance with varying $\lambda$’s can be found in Appendix A.4.1.
Table 1: Comparison of accuracy on MultiBench datasets (max over $\lambda$, averaged over 10 seeds).
@@
Our method (IndiSeek) 56.81% 73.87% 60.43% 53.91% 65.80%
InfoDisen 55.74% 73.17% 57.67% 53.41% 65.58%
FactorizedCL 55.92% 73.12% 57.61% 53.70% 65.69%
CLIP (baseline) 56.11% 72.49% 56.81% 53.77% 64.60%
To further evaluate the learned task-agnostic representations, we also investigate the multi-task performance of each method on the MIMIC dataset. We consider the tasks of predicting ICD-9 codes for multiple groups in the MIMIC dataset in MultiBench (Liang et al.,, 2021). We compare the performance of all three disentangled learning methods together with CLIP baseline in Table 2, in which we calculate the average accuracy across tasks of predicting the last three ICD-9 codes, and present results in a subset of ICD-9 codes. Complete results for other tasks are deferred to the appendix.
Table 2: Multi-task performance on MIMIC dataset (ICD-9 groups 17–19).
@@
IndiSeek 62.26% 60.92% 69.56% 64.25%
FactorizedCL 61.88% 60.53% 69.43% 63.95%
InfoDisen 61.90% 60.57% 69.50% 63.99%
CLIP 61.33% 60.80% 69.33% 63.82%
@@ 6 Extensions and discussion
  • With task-specific augmentations. When a downstream task is known in advance, domain knowledge can be used to design within-modality data augmentations. If the augmentations are ideal in the sense that the task-invariant orbit of each observation is fully covered (Liang et al.,, 2023), one may first perform within-modality contrastive learning. The resulting representations retain only task-relevant information. IndiSeek can then be applied to further decompose these task-relevant features into shared and modality-specific components.
@@
Our paper also leaves an open question regarding how to select the Lagrange multiplier $\lambda$ in (6) in a principled manner. As the CLUB upper bound and the squared reconstruction loss are at different scales, the optimal $\lambda$ may itself be scale-dependent and vary across datasets. Identifying the optimal monotone transformation of the squared reconstruction loss remains an interesting future direction to enhance empirical performance.
@@
In the literature, there are other alternatives that learn both shared and modality-specific features in an information-theoretic manner. For example, Dufumier et al., (2024) adopts infoNCE as a proxy for mutual information. However, the framework in Dufumier et al., (2024) cannot ensure disentanglement between shared and modality-specific features, leading to potential redundancy in learned representations. In a concurrent work, Shi et al., (2025) adopts variational bounds for mutual information and entropy regularization as objectives to learn disentangled features. However, similar to Liang et al., (2023) and Dufumier et al., (2024), the objectives are motivated by some task labels $Y$ and hence differ from our goal of learning task-agnostic features.
@@ Acknowledgements
C.M. was partially supported by the National Science Foundation via grant DMS-2311127. Z.M. was partially supported by the National Institutes of Health via grant U01CA294514 and the National Science Foundation via grants DMS-2345215 and DMS-2245575.
@@ A.1 Details on metrics
In simulated experiments, to determine which coordinates are extracted as shared or modality-specific features in $X_{1}$, we consider model-free feature importance based on feature masking related to Robnik-Šikonja and Kononenko, (2008); Zeiler and Fergus, (2014); Li et al., (2016). Concretely, given a black-box architecture $\psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{p}$, to understand which coordinates in each $x\in\mathbb{R}^{d}$ are effective in the output $\psi(x)$, we use leave-one-coordinate-out $x_{-j}$ as inputs for $j\in[d]$ and quantify the difference between and original output and output with masked input $\|\psi(x)-\psi(x_{-j})\|^{2}$. In implementation, for each coordinate $j\in[d]$, we draw $M=1000$ random vectors $\{x_{i}\}_{i\in[M]}$ from uniform distributions on $10,10]^{d}$ and calculate the average
@@
Recall that we obtain disentangled representations for RNA and ADT data, $(C_{\rm RNA},C_{\rm ADT})$ and $(Z_{\rm RNA}$, $Z_{\rm ADT})$, using the training set. To this end, for each cell $c$, we find its 10 nearest neighbors measured by Euclidean distance in $Z_{\rm RNA}$ and $Z_{\rm ADT}$, respectively, and compare the proportions of neighbors with the same level-2 annotations as $i$ in the two modalities, denoted by $\beta_{\rm RNA}(c)$ and $\beta_{\rm ADT}(c)$, respectively. We summarize the comparison with a score $\theta(c)=\log(\beta_{\rm RNA}(c)/\beta_{\rm ADT}(c))$ that is monotone increasing with respect to the importance of RNA-specific information, and rank $27$ cell types based on this metric. We use the RNA weights based on weighted k-NN obtained in Hao et al., (2021) as the benchmark and compare our ranks $(R_{i})_{i\in[L]}$ for cell types with the RNA-weights-based ranks $(R^{*}_{i})_{i\in[L]}$ with $L=27$. We adopt Spearman’s $\rho$ correlation as the metric, where for any two vectors of ranks as permutations of $[L]$,
@@ A.2 Numerical simulations
For Setting 1, given shared features are generated by sine and cubic functions, we can see from Figure 6 and 7 that both Factorized CL and IndiSeek are capable of extracting modality-specific features that are independent of shared ones when $\lambda=0.01$, while IndiSeek tend to be more robust to the choice of $\lambda$ and preserves the desired performance when $\lambda=0.1$. However, in contrast, InfoDisen, due to the limitation of orthogonal loss in handling nonlinear dependence, fails to disentangle modal-specific features from shared ones regardless of the value of $\lambda$.
@@ A.3 CITE-seq dataset implementation details
We consider the bone marrow CITE-seq dataset from Stuart et al., (2019) which consists of measurements in two modalities on $30672$ individual cells: transcriptome (RNA) and $25$ cell-surface proteins (ADT). We randomly split the dataset into a training set of size $15000$ and a test set of size $15672$. We follow the pre-processing in Hao et al., (2021), [^7] https://satijalab.org/seurat/articles/weighted_nearest_neighbor_analysis where we conduct scaling and centering based on moments estimated from the training set only.
@@ A.4 MultiBench datasets
We evaluate the performance of our method on MultiBench datasets Liang et al., (2021), including video datasets (MOSI (Zadeh et al.,, 2016), MOSEI (Zadeh et al.,, 2018), UR-FUNNY (Hasan et al.,, 2019), MUStARD (Castro et al.,, 2019)) with modalities of text, image, and audio, and the MIMIC dataset (Johnson et al.,, 2016). For video datasets, following Liang et al., (2023); Wang et al., 2024a , we focus solely on vision and text modalities, excluding audio data in these examples. More concretely,
@@
  • MOSI Zadeh et al., (2016): $2199$ YouTube clips with vision and text modalities. Task: binary sentiment classification.
@@
  • MOSEI (Zadeh et al.,, 2018): $23000$ monologue videos with vision and text modalities. Task: binary sentiment classification.
@@
  • UR-FUNNY (Hasan et al.,, 2019): TED talk videos with vision and text modalities. Task: binary humor detection.
@@
  • MUStARD (Castro et al.,, 2019): $690$ TV show clips with vision and text modalities. Task: binary sarcasm detection.
@@
  • MIMIC (Johnson et al.,, 2016): over $36000$ ICU patient records with time-series vitals and tabular data. Task: Following Wang et al., 2024a , we aim at predicting ICD-9 group $7$ membership.
@@
We follow the processing pipeline in Liang et al., (2023) and adopt exactly the same data splitting and feature pre-extraction. We use a smaller Transformer architecture with $2$ heads, $2$ layers, and the intermediate layers with a width $128$.
Table 3: MOSEI results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 72.45% 72.45% 72.50% 73.16% 73.33% 73.65% 73.87% 73.87%
FactorizedCL 72.47% 72.54% 72.64% 73.12% 72.82% 72.58% 72.61% 73.12%
InfoDisen 72.94% 73.17% 72.72% 72.75% 72.87% 72.90% 72.99% 73.17%
CLIP 72.49% 72.49% 72.49% 72.49% 72.49% 72.49% 72.49% 72.49%
@@
Table 4: MOSI results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 56.11% 56.31% 55.93% 56.22% 56.59% 56.81% 56.30% 56.81%
FactorizedCL 55.79% 55.69% 55.92% 54.50% 54.60% 55.38% 54.83% 55.92%
InfoDisen 54.61% 55.07% 54.81% 55.05% 55.04% 55.74% 55.74% 55.74%
CLIP 56.11% 56.11% 56.11% 56.11% 56.11% 56.11% 56.11% 56.11%
@@
Table 5: MUStARD results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 52.54% 52.17% 52.97% 53.91% 53.84% 53.41% 52.46% 53.91%
FactorizedCL 53.70% 53.51% 53.12% 52.10% 51.85% 53.04% 52.79% 53.70%
InfoDisen 50.29% 49.71% 49.64% 52.83% 53.41% 53.33% 53.33% 53.41%
CLIP 53.77% 53.77% 53.77% 53.77% 53.77% 53.77% 53.77% 53.77%
@@
Table 6: UR-FUNNY results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 56.81% 56.82% 57.07% 59.79% 59.79% 60.43% 60.20% 60.43%
FactorizedCL 56.87% 57.61% 57.59% 56.75% 57.39% 57.04% 57.17% 57.61%
InfoDisen 56.30% 55.99% 56.10% 56.44% 56.41% 57.67% 56.89% 57.67%
CLIP 56.81% 56.81% 56.81% 56.81% 56.81% 56.81% 56.81% 56.81%
@@
Table 7: MIMIC results with varying $\lambda$ (averaged over 10 seeds).
$\lambda$ 0.001 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 64.64% 65.20% 65.44% 65.64% 65.68% 65.65% 65.80% 65.80%
FactorizedCL 64.57% 65.15% 65.09% 65.58% 65.69% 65.52% 65.58% 65.69%
InfoDisen 64.65% 64.80% 64.73% 64.75% 65.02% 65.58% 65.38% 65.58%
CLIP 64.60% 64.60% 64.60% 64.60% 64.60% 64.60% 64.60% 64.60%

### v3

@@ 5 Experiments with MultiBench datasets
In addition to the application in single-cell multi-omics, we also evaluate the performance of IndiSeek on MultiBench datasets (liang2021multibench), which include 4 video datasets (MOSI (zadeh2016mosi), MOSEI (zadeh2018multimodal), UR-FUNNY (hasan2019ur), MUStARD (castro2019towards)) and Medical Information Mart for Intensive Care III (MIMIC) dataset (johnson2016mimic). We follow the processing pipeline in liang2023factorized and adopt exactly the same data splitting and feature pre-extraction, in which we focus on the task of binary classification on whether the patient fits any ICD-9 code in group 7 for the MIMIC dataset. We use a smaller Transformer architecture with $2$ heads, $2$ layers, and the intermediate layers with a width $128$. For all methods in comparison, we train the architecture for $2000$ epochs with the same training parameters. We note that our focus here is to benchmark the qualities of representations learned with task-agnostic objective functions. Therefore, for fair comparison along this direction, we do not involve task-related within- or cross-modality data augmentations and hence do not involve self-supervised learning terms in objective functions of all methods in comparison. In contrast, to achieve the optimal performance in a specific task on a particular benchmark dataset, appropriate data augmentation informed by the task and domain knowledge is often helpful.
Following liang2023factorized and wang2024information, we focus on the linear probing accuracy for each method with learned representations $(C_{1},Z_{1},C_{2},Z_{2})$ as features, i.e., we concatenate shared and modality-specific features for classification. Here we also compare with the CLIP baseline, where only the shared representations $(C_{1},C_{2})$ are used for linear probing accuracy evaluation. Following the same training procedure as for the CITE-seq dataset, we report the results with three methods in Table 1. For each method, we vary $\lambda$ in $\{10^{j}:j=-3,-2,-1,0,1,2,3\}$ and for each $\lambda$, we run each method with $10$ random seeds. Table 1 presents the averaged accuracy for the best choice of $\lambda$. More details on the performance with varying $\lambda$’s can be found in Appendix A.4.1.
Ablation studies with varying output dimension.
To investigate the robustness of IndiSeek across different output dimension capacities, we conduct experiments on the MOSI dataset with output dimensions $d\in\{20,60,100,140,180\}$. For each output dimension, we perform a sweep over $\lambda$ and report the maximum accuracy achieved across all $\lambda$ values, averaged over 10 random seeds. Table 2 presents the results of this analysis. IndiSeek consistently achieves the best performance across all output dimensions, demonstrating its robustness to this hyperparameter choice.
Table 1: Comparison of accuracy on multimodal datasets (max over $\lambda$, averaged over 10 seeds). All values are percentages.
@@
IndiSeek 70.03(1.39) 75.47(0.13) 63.79(0.39) 57.46(1.04) 65.99(0.31)
FactorizedCL 67.11(0.34) 74.74(0.04) 58.36(0.25) 56.45(1.16) 65.69(0.11)
InfoDisen 67.52(0.62) 74.73(0.08) 58.08(0.50) 56.16(0.92) 65.47(0.23)
CLIP (baseline) 67.61(0.66) 74.70(0.05) 58.32(0.68) 55.36(1.12) 64.56(0.29)
Table 2: MOSI results with varying output dimensions (max over $\lambda$, averaged over 10 seeds). All values are percentages.
Method outdim=20 outdim=60 outdim=100 outdim=140 outdim=180
IndiSeek 69.68(0.51) 70.03(0.80) 69.56(0.57) 70.20(0.65) 68.80(1.15)
FactorizedCL 69.49(0.26) 66.91(1.04) 67.52(0.95) 68.28(0.67) 67.08(0.29)
InfoDisen 68.57(0.91) 65.95(1.59) 65.17(0.68) 66.90(0.33) 66.62(0.68)
CLIP (baseline) 69.46(0.45) 67.03(0.79) 66.53(0.68) 68.54(0.94) 67.20(1.04)
Multi-task performance.
To further evaluate the learned task-agnostic representations, we also investigate the multi-task performance of each method on the MIMIC dataset. We consider the tasks of predicting ICD-9 codes for multiple groups in the MIMIC dataset in MultiBench (liang2021multibench). We compare the performance of all three disentangled learning methods together with CLIP baseline in Table 3, in which we calculate the average accuracy across tasks of predicting the last three ICD-9 codes, and present results in a subset of ICD-9 codes. Complete results for other tasks are deferred to the appendix.
Table 3: Multi-task performance on MIMIC dataset (ICD-9 groups 17–19). All values are percentages.
@@
IndiSeek 62.58(0.17) 60.99(0.19) 69.71(0.11) 64.43
FactorizedCL 61.83(0.20) 60.36(0.14) 69.30(0.12) 63.83
InfoDisen 61.72(0.14) 60.40(0.34) 69.46(0.14) 63.86
CLIP (baseline) 61.50(0.14) 60.71(0.09) 69.25(0.16) 63.82
@@ 6 Extensions and discussion
  • With task-specific augmentations. When a downstream task is known in advance, domain knowledge can be used to design within-modality data augmentations. If the augmentations are ideal in the sense that the task-invariant orbit of each observation is fully covered (liang2023factorized), one may first perform within-modality contrastive learning. The resulting representations retain only task-relevant information. IndiSeek can then be applied to further decompose these task-relevant features into shared and modality-specific components.
@@
The selection of the Lagrange multiplier $\lambda$ in (6) is important for the implementation of IndiSeek. With the rescaled reconstruction loss, the CLUB upper bound and the reconstruction term are now both scale-free, and our experimental results have suggested a candidate range of $\lambda$’s between $10^{-1}$ to $10^{1}$. Identifying the precise optimal choice of $\lambda$ remains an interesting future direction to enhance empirical performance and theoretical understanding.
@@
In the literature, there are other alternatives that learn both shared and modality-specific features in an information-theoretic manner. For example, dufumier2024align adopts infoNCE as a proxy for mutual information. However, the framework in dufumier2024align cannot ensure disentanglement between shared and modality-specific features, leading to potential redundancy in learned representations. In a concurrent work, shi2025towards adopts variational bounds for mutual information and entropy regularization as objectives to learn disentangled features. However, similar to liang2023factorized and dufumier2024align, the objectives are motivated by some task labels $Y$ and hence differ from our goal of learning task-agnostic features.
@@ Acknowledgements
C.M. was partially supported by the National Science Foundation via grant DMS-2311127 and the CAREER Award DMS-2443867. Z.M. was partially supported by the National Institutes of Health via grant U01CA294514 and the National Science Foundation via grants DMS-2345215 and DMS-2245575.
@@ A.1 Implementation details and metrics
In simulated experiments, to determine which coordinates are extracted as shared or modality-specific features in $X_{1}$, we consider model-free feature importance based on feature masking related to robnik2008explaining; zeiler2014visualizing; li2016understanding. Concretely, given a black-box architecture $\psi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{p}$, to understand which coordinates in each $x\in\mathbb{R}^{d}$ are effective in the output $\psi(x)$, we use leave-one-coordinate-out $x_{-j}$ as inputs for $j\in[d]$ and quantify the difference between and original output and output with masked input $\|\psi(x)-\psi(x_{-j})\|^{2}$. In implementation, for each coordinate $j\in[d]$, we draw $M=1000$ random vectors $\{x_{i}\}_{i\in[M]}$ from uniform distributions on $10,10]^{d}$ and calculate the average
@@
Recall that we obtain disentangled representations for RNA and ADT data, $(C_{\rm RNA},C_{\rm ADT})$ and $(Z_{\rm RNA}$, $Z_{\rm ADT})$, using the training set. To this end, for each cell $c$, we find its 10 nearest neighbors measured by Euclidean distance in $Z_{\rm RNA}$ and $Z_{\rm ADT}$, respectively, and compare the proportions of neighbors with the same level-2 annotations as $i$ in the two modalities, denoted by $\beta_{\rm RNA}(c)$ and $\beta_{\rm ADT}(c)$, respectively. We summarize the comparison with a score $\theta(c)=\log(\beta_{\rm RNA}(c)/\beta_{\rm ADT}(c))$ that is monotone increasing with respect to the importance of RNA-specific information, and rank $27$ cell types based on this metric. We use the RNA weights based on weighted k-NN obtained in hao2021integrated as the benchmark and compare our ranks $(R_{i})_{i\in[L]}$ for cell types with the RNA-weights-based ranks $(R^{*}_{i})_{i\in[L]}$ with $L=27$. We adopt Spearman’s $\rho$ correlation as the metric, where for any two vectors of ranks as permutations of $[L]$,
@@
Implementation of CLUB loss.
We follow the implementation of CLUB in cheng2020club. As CLUB loss is a variational upper bound for mutual information, we use an auxiliary conditional density $q_{\theta}(z|c)$ as an approximation for $p(c|z)$, where $q_{\theta}(z|c)$ lies in the Gaussian family, and we optimize the moments of the Gaussian distribution using 5-layer MLPs in each epoch. The inner optimization is run for 5 epochs at each epoch to optimize representations, and the implementation is the same for both IndiSeek (we use cross-modal shared features) and FactorCL for fair comparison.
For FactorizedCL, liang2021multibench also proposes a self-supervised variant of Factorized CL by leveraging task-relevant augmentations, more discussions on which are deferred to Section 6. In this paper, we adapt the unsupervised version of Factorized CL, i.e., we use CLUBInfoNCECritic between $Z_{1}$ and $Z_{2}$ in liang2023factorizedgithub as the disentanglement objective, and adopt the implementation of CLUB loss in cheng2020club in both IndiSeek  and Factorized CL for fair comparison.
@@ A.2 Numerical simulations
For Setting 1, given shared features are generated by sine and cubic functions, we can see from Figure 6 and 7 that both Factorized CL and IndiSeek are capable of extracting modality-specific features that are independent of shared ones when $\lambda=0.01$, while IndiSeek tend to be more robust to the choice of $\lambda$ and preserves the desired performance when $\lambda=0.1$. However, in contrast, InfoDisen, due to the limitation of orthogonal loss in handling nonlinear dependence, fails to disentangle modality-specific features from shared ones regardless of the value of $\lambda$.
@@ A.3 CITE-seq dataset implementation details
We consider the bone marrow CITE-seq dataset from stuart2019comprehensive which consists of measurements in two modalities on $30672$ individual cells: transcriptome (RNA) and $25$ cell-surface proteins (ADT). We randomly split the dataset into a training set of size $15000$ and a test set of size $15672$. We follow the preprocessing in https://satijalab.org/seurat/articles/weighted_nearest_neighbor_analysis, where we normalize ADT data with centering to produce a 24-dimensional input based on moments estimated from the training set only, and normalize the extremely high-dimensional RNA data and extract the first 200 principal components as the inputs for CLIP.
@@
In Figure 18, Figure 19, and Figure 20, we also visualize the learned representations $(C_{i},Z_{i})$ by all three methods with $\lambda=1.0$ for each modality. In the left panels, the representations are colored by level-1 cell types, and in the right panels, to ease visualization, we plot a subset of representations with level-2 cell types that are among the top-$40\%$ in the benchmark weights in the corresponding modality.
To quantify the separation of clusters, we adopt two metrics: (1) the Adjusted Rand Index (ARI) (hubert1985comparing), which measures agreement between $k$-means clustering on the learned representations and the ground-truth labels, with values ranging from 1-1 to $1$ where higher values indicate better cluster recovery; and (2) the Silhouette Score (rousseeuw1987silhouettes), which quantifies cluster compactness and separation using the ground-truth labels, with values in [1,1][-1,1] where higher scores indicate better-defined clusters.
Figure 18: UMAP of IndiSeek-learned representations $(C_{i},Z_{i})$ (colored by cell types).
Figure 19: UMAP of FactorizedCL-learned representations $(C_{i},Z_{i})$ by IndiSeek (colored by cell types).
Figure 20: UMAP of InfoDisen-learned representations $(C_{i},Z_{i})$ by IndiSeek (colored by cell types).
@@ A.4 MultiBench datasets
We evaluate the performance of our method on MultiBench datasets liang2021multibench, including video datasets (MOSI (zadeh2016mosi), MOSEI (zadeh2018multimodal), UR-FUNNY (hasan2019ur), MUStARD (castro2019towards)) with modalities of text, image, and audio, and the MIMIC dataset (johnson2016mimic). For video datasets, following liang2023factorized; wang2024information, we focus solely on vision and text modalities, excluding audio data in these examples. More concretely,
@@
  • MOSI zadeh2016mosi: $2199$ YouTube clips with vision and text modalities. Task: binary sentiment classification.
@@
  • MOSEI (zadeh2018multimodal): $23000$ monologue videos with vision and text modalities. Task: binary sentiment classification.
@@
  • UR-FUNNY (hasan2019ur): TED talk videos with vision and text modalities. Task: binary humor detection.
@@
  • MUStARD (castro2019towards): $690$ TV show clips with vision and text modalities. Task: binary sarcasm detection.
@@
  • MIMIC (johnson2016mimic): over $36000$ ICU patient records with time-series vitals and tabular data. Task: Following wang2024information, we aim at predicting ICD-9 group $7$ membership.
@@
We follow the processing pipeline in liang2023factorized and adopt exactly the same data splitting and feature pre-extraction. We use a smaller Transformer architecture with $2$ heads, $2$ layers, and the intermediate layers with a width $128$.
@@
Table 4: MOSI results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
$\lambda$ 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 67.46(0.74) 69.48(0.86) 70.03(1.39) 66.43(0.34) 65.95(0.45) 65.90(0.89) 70.03(1.39)
FactorizedCL 66.76(0.58) 65.07(0.92) 66.12(0.67) 64.26(1.07) 65.07(0.58) 65.12(0.89) 67.11(0.34)
InfoDisen 65.57(0.84) 65.31(1.26) 65.83(1.00) 65.22(1.08) 64.84(1.26) 67.52(0.62) 67.52(0.62)
CLIP 67.61(0.66) 67.61(0.66) 67.61(0.66) 67.61(0.66) 67.61(0.66) 67.61(0.66) 67.61(0.66)
@@
Table 5: MOSEI results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
$\lambda$ 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 74.70(0.06) 74.71(0.05) 75.47(0.13) 75.29(0.09) 75.19(0.11) 75.03(0.11) 75.47(0.13)
FactorizedCL 74.74(0.04) 74.74(0.05) 74.65(0.04) 74.69(0.07) 74.65(0.04) 74.60(0.06) 74.74(0.04)
InfoDisen 74.65(0.06) 74.61(0.03) 74.58(0.04) 74.70(0.03) 74.73(0.08) 74.56(0.04) 74.73(0.08)
CLIP 74.70(0.05) 74.70(0.05) 74.70(0.05) 74.70(0.05) 74.70(0.05) 74.70(0.05) 74.70(0.05)
@@
Table 6: UR-FUNNY results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
$\lambda$ 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 58.85(1.08) 62.40(0.82) 63.08(0.44) 63.79(0.39) 63.12(0.55) 63.71(0.62) 63.79(0.39)
FactorizedCL 58.36(0.25) 57.81(0.41) 58.22(0.38) 57.43(0.64) 57.13(0.24) 57.02(0.32) 58.36(0.25)
InfoDisen 56.09(0.33) 56.14(0.98) 56.28(0.76) 57.52(0.57) 58.08(0.50) 57.73(0.70) 58.08(0.50)
CLIP 58.32(0.68) 58.32(0.68) 58.32(0.68) 58.32(0.68) 58.32(0.68) 58.32(0.68) 58.32(0.68)
@@
Table 7: MUStARD results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
$\lambda$ 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 55.51(1.28) 57.46(1.04) 55.72(1.95) 55.07(1.06) 53.48(1.42) 52.25(1.17) 57.46(1.04)
FactorizedCL 53.91(0.57) 54.71(1.21) 54.86(1.10) 54.64(0.70) 56.45(1.16) 55.07(0.91) 56.45(1.16)
InfoDisen 54.13(0.83) 53.41(1.00) 54.42(0.97) 54.57(0.94) 56.16(0.92) 55.58(1.27) 56.16(0.92)
CLIP 55.36(1.12) 55.36(1.12) 55.36(1.12) 55.36(1.12) 55.36(1.12) 55.36(1.12) 55.36(1.12)
Table 8: MIMIC results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
$\lambda$ 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 65.52(0.22) 65.79(0.17) 65.82(0.24) 65.89(0.24) 65.96(0.28) 65.99(0.31) 65.99(0.31)
FactorizedCL 65.34(0.29) 65.11(0.31) 65.69(0.11) 65.33(0.30) 64.92(0.35) 65.24(0.39) 65.69(0.11)
InfoDisen 64.89(0.30) 64.80(0.31) 65.10(0.25) 64.96(0.38) 65.47(0.23) 65.32(0.24) 65.47(0.23)
CLIP 64.56(0.29) 64.56(0.29) 64.56(0.29) 64.56(0.29) 64.56(0.29) 64.56(0.29) 64.56(0.29)
Table 9: Comparison of training time on multimodal datasets (max over $\lambda$, averaged over 10 seeds). All values are in seconds ($\times 10^{2}$).
Method MOSI MOSEI UR-FUNNY MUStARD MIMIC
IndiSeek 7.11(0.09) 81.43(0.83) 40.75(0.29) 2.63(0.04) 132.74(1.76)
FactorizedCL 7.35(0.09) 83.27(0.87) 41.97(0.31) 2.72(0.04) 135.93(1.84)
InfoDisen 7.18(0.09) 81.32(0.85) 40.70(0.30) 2.65(0.04) 132.60(1.98)